Is this statement true or false
The ordered pair (0,8) is the solution of the following system of equations.
-8x-7y=-56
5x+9y=72

Answer by nabla(475)

The ordered pair will be in form (x,y). So you have x=0, y=8. Put those values into the left side of each equation to see if you result in the respective right side.
-8(0)-7(8)=0-56=-56.
5(0)+9(8)=0+72=72.
These sums of the left equal the numbers on the right. Thus, the ordered pair is a solution to the system of equations.